OpenStudy (anonymous):

The numberof. Subscribers yet o a website after t years is shown by the equation below. Y=50(1.75)^t. Which conclusion is correct about the number of subscribers to the website. A. It increased by 75% each year B. It decreased by 75% each year C it increased by 50% every year D. It decreased by 50% every year

OpenStudy (anonymous):

@hero

hero (hero):

@kristinak, which choice do you believe is correct and why?

hero (hero):

You might want to insert values of t to help you understand: Y=50(1.75)^0 Y=50(1.75)^1 Y=50(1.75)^2

OpenStudy (anonymous):

I believe it is A based on the equation

OpenStudy (anonymous):

@Hero

hero (hero):

Can you explain why A is correct?

hero (hero):

Being able to explain why it is correct increases your understanding of the concept.

OpenStudy (anonymous):

it is correct because if in that equation 1.75 represents the 75% in that form and the exponent gives me that idea that number is going to keep on increasing

hero (hero):

The equation you posted was in the form \(y = a^{bt}\) where b is the growth factor if b > 1 resulting in an increasing function decay factor if b 0 < x < 1 resulting in a decreasing function. Now since in this case b > 1, we have the form b = (1 + r) where r is the rate in decimal form. In this case r = .75 and b = (1 + .75) so the function will definitely increase. And since r = .75, it will increase by 75% incrementally.